OpenStudy (anonymous):

Find the product 7n^4 * 2n^5

Nnesha (nnesha):

when you mulitplyc same bases you should add their exponents exxample \[\huge\rm x^m \times x^n = x^{m+n}\]

OpenStudy (anonymous):

i mean 14^9

Nnesha (nnesha):

9? to the 9 power ? multiply coefficient (7 times 2)(n^4 times x^5)

Nnesha (nnesha):

14 is not to the 9 power the variable is to the 9 poawer

OpenStudy (anonymous):

i am still so confused

OpenStudy (anonymous):

so like 14^20

Nnesha (nnesha):

nope 9 is right bec you have to add exponents so n^9 not 14^9

OpenStudy (anonymous):

okay so it would be 14n^9

Nnesha (nnesha):

\[\huge\rm (7 \times 2)(n^4 \times n^5)=(7 \times 2) n^{4+5}\]

Nnesha (nnesha):

yes that's right

OpenStudy (anonymous):

5xy^5 * -3x^4y^3 how would that one be set up to find the product

Nnesha (nnesha):

* <--is that a multiplicationz sign ?

OpenStudy (anonymous):

yea

Nnesha (nnesha):

that's same like we just did multiply coefficients and add exponents of variables

Nnesha (nnesha):

x is same as x to the one power x=x^1

OpenStudy (anonymous):

so it would be like -15^xy13? or -15x^5y8

OpenStudy (anonymous):

-15x^5y^8 i wrote that wrong

Nnesha (nnesha):

add coefficient of same variables you can't add 3 exponent of y with 5 exponent of x x^{1+4}

Nnesha (nnesha):

so which one is correct ?

OpenStudy (anonymous):

so it would be -15x^5y^8

OpenStudy (anonymous):

thank you it was right

Nnesha (nnesha):

np :=)